Solve by completing the square: x^2- 4x-11=0
Answer by edjones(8007)

x^2-4x-11=0
x^2-4x =11
x^2-4x+4=11+4
(x-2)^2=15
x-2=+-sqrt(15)
x=2+sqrt(15), x=2-sqrt(15)
.
Ed